# Approximating Sparse Matrices and their Functions using Matrix-vector products

### v2

@@ Abstract
Matrix function, banded matrix, sparse matrix, matrix-vector products, compressed sensing
MSC:
[2020] 65F50, 65F55, 65F60
journal: Applied and Computational Harmonic Analysis
@@ 1 Introduction
The evaluation of matrix functions is an important task arising frequently in many areas of scientific computing. They appear in the numerical solution of partial differential equations [24, 32, 36], network analysis [4, 26], electronic structure calculations [33, 49] and statistical learning [17, 58], among others. In some problems, the matrix of interest has some underlying structure such as bandedness or sparsity. For example, in graphs with community structure one requires functions of sparse matrices [52] and for electronic structure calculations one computes functions of banded matrices [33, 49]. In this paper, we consider the problem of approximating $f(\bm{A})$ through matrix-vector products $\bm{x}\mapsto\bm{Ax}$ only, assuming $\bm{A}$ is sparse or banded.
@@
There are many other equivalent ways of defining $f(\bm{A})$, see for example [42].
@@
When the matrix dimension is large, explicitly computing $f(\bm{A})$ using standard algorithms [34, 42] becomes prohibitive in practice, usually requiring $O(n^{3})$ operations. Moreover, the matrix $\bm{A}$ may also be too large to store explicitly. In this scenario, $\bm{A}$ is often only available through matrix-vector products and $f(\bm{A})\bm{b}$  [^1] We note that the matrix-vector product $f(\bm{A})\bm{b}$ is never computed explicitly by finding $f(\bm{A})$ and then multiplying $\bm{b}$ to $f(\bm{A})$, but with matrix-vector product with $\bm{A}$ only using, for example, the Krylov methods. is used to infer information about $f(\bm{A})$ [37, 38]. In this paper, we show that the whole matrix $f(\bm{A})$ can still be approximated in a sparse format using $\ll n$ matrix-vector products with $\bm{A}$ if $\bm{A}$ is sparse and $f$ is smooth. We exploit the decay bounds for the entries of $f(\bm{A})$ [8] to recover its large entries using matrix-vector products;see also [3]. We primarily study functions of sparse matrices with general sparsity pattern and matrices with few large entries in each row/column (See Section 2), but also consider functions of banded matrices and matrices with off-diagonal decay as a special case (See Section 3). Our work is particularly useful if $f(\bm{A})$ is sparse or has only a few entries that are large in each row/column or if $\bm{A}$ is banded with bandwidth $k\ll n$.
@@
Let $\bm{A}$ be a diagonalizable, sparse matrix with some sparsity pattern and $f$ (either $f:\mathbb{R}\rightarrow\mathbb{R}$ or $f:\mathbb{C}\rightarrow\mathbb{C}$) a function that is analytic in a region containing the spectrum of $\bm{A}$. Depending on the sparsity pattern of $\bm{A}$ and the smoothness of $f$, some entries of $f(\bm{A})$ can be shown to be exponentially smaller than others [8] and we can view them as noise. For a sparse matrix with a general sparsity pattern, it can be difficult to know the support (location of the large entries) of $f(\bm{A})$ in advance, for example, if the sparsity pattern of $\bm{A}$ is unknown. However, we can write $f(\bm{A})=\widehat{f(\bm{A})}+\bm{E}$ with the matrix $\bm{E}$ having only small entries in magnitude and compute $\widehat{f(\bm{A})}$ by viewing $\widehat{f(\bm{A})}$ as signals and $\bm{E}$ as noise using compressed sensing row by row or column by column. Compressed sensing is a technique in signal processing for efficiently reconstructing a signal which is sparse in some domain by finding solutions to underdetermined linear systems [13, 15, 23]. As a special case (by taking $f(\bm{A})=\bm{A}$), our algorithm can be used for recovering a sparse matrix $\bm{A}$ itself, where $\bm{A}$ is accessed only through matrix-vector multiplications $\bm{x}\mapsto\bm{A}\bm{x}$. The details will be discussed in Section 2.
@@
Let $\bm{A}$ now be a banded matrix. Then $f(\bm{A})$ can usually be well approximated by a banded matrix [22]. There are theoretical results that confirm this property [7], which show that the entries of $f(\bm{A})$ decay exponentially away from the main diagonal. The observation is that $\left|[f(\bm{A})]_{ij}\right|\ll\left|[f(\bm{A})]_{ii}\right|,\left|[f(\bm{A})]_{jj}\right|$ for $|i-j|\gg 1$. Therefore, we can dismiss the exponentially small entries and find a matrix that recovers the large entries of $f(\bm{A})$, i.e., find a matrix $\widehat{f(\bm{A})}$ such that the difference $f(\bm{A})-\widehat{f(\bm{A})}$ only contains small entries in magnitude. We can show that the large entries of $f(\bm{A})$ can be captured through matrix-vector products with the following simple deterministic matrix
@@
where $s<n$, $\bm{I}_{s}$ is the $s\times s$ identity matrix and $\bm{I}_{s}(:,1:\{n/s\}\cdot s)$ are the first $\{n/s\}\cdot s$ columns of $\bm{I}_{s}$, which ensures that $\bm{I}_{n}^{(s)}$ has exactly $n$ rows. Here $\{x\}$ is the fractional part of $x$. This is unlike the sparse case where we now have significant information on the support of $f(\bm{A})$. This set of vectors $\bm{I}_{n}^{(s)}$ are identical to the probing vectors proposed in [31] for estimating functions of banded matrices and [16, 19] for estimating sparse, banded Jacobian matrices. If we let $s$ be proportional to the bandwidth of $\bm{A}$, the matrix-vector products $f(\bm{A})\bm{I}_{n}^{(s)}$ capture the large entries near the main diagonal, giving us a good approximation for $f(\bm{A})$. We discuss the details in Section 3.
Both cases require matrix-vector products with $f(\bm{A})$, i.e., $\bm{x}\mapsto f(\bm{A})\bm{x}$. Popular methods for this include the standard (polynomial) Krylov method [42, 56], the rational Krylov method [37] and contour integration [39]. In this paper, we discuss three functions: the exponential $e^{\bm{A}}$, the square root function $\sqrt{\bm{A}}$ and the logarithm $\log(\bm{A})$. For the exponential we use the polynomial Krylov method, and for the square root function and the logarithm, we use contour integration as in [39, method 2] as both functions have a branch cut on the negative real axis $(-\infty,0]$, making them difficult to approximate using polynomials if $\bm{A}$ has eigenvalues close to 00. If $f$ has poles in the region containing the spectrum of $\bm{A}$ or if $f$ cannot be well-approximated by a low-degree polynomial, then rational Krylov methods can also be used. For Krylov methods, block versions can also be used [29], which compute $f(\bm{A})\bm{B}\in\mathbb{R}^{n\times s}$ at once instead of column by column. We do not discuss the details of computing the matrix-vector product $f(\bm{A})\bm{b}$ further as they are not the focus of this paper.
@@ 1.1 Existing methods
First, Benzi and Razouk [8] describe an algorithm that exploits the decay property by first finding a good polynomial approximation $p$ to $f$, for example using Chebyshev interpolation, and computing $p(\bm{A})$ while using a dropping strategy to keep the matrix as sparse as possible. However, this method can quickly become infeasible if $f$ cannot be well approximated by a low-degree polynomial or if sparse data structures and arithmetic is not available. Frommer, Schimmel and Schweitzer [31] use probing methods to approximate matrix functions $f(\bm{A})$ and its related quantities, for instance the trace of $f(\bm{A})$. The method finds probing vectors to estimate the entries of $f(\bm{A})$ by partitioning the vertices of the directed graph $G(\bm{A})$ associated with the sparse matrix $\bm{A}$ using graph coloring and the sparsity pattern of $\bm{A}$. Our Section 3 on banded matrices can be seen as a special case of [31], but as banded matrices arise naturally in applications and come with strong theoretical results, we make the sensing matrix explicit and study them in detail. The strategy proposed in [31] costs $\mathcal{O}(\Delta^{d}n)$ to obtain the graph coloring where $\Delta$ is the maximal degree of $G(\bm{A})$ and $d$ is a distance parameter. In the special case when $\bm{A}$ is a banded matrix the graph coloring can be obtained with $\mathcal{O}(n)$ cost. Our work and [31] satisfies a similar theoretical accuracy bound up to a low-degree polynomial factor in $n$ (cf. Section 3 of [31] and Theorem 1 in our work), but we use compressed sensing to estimate the large entries of each row/column. Cortinovis, Kressner and Massei [18] use a divide-and-conquer approach to approximate matrix functions where the matrix is either banded, hierarchically semiseparable or has a related structure. The divide-and-conquer method is based on (rational) Krylov subspace methods for performing low-rank updates of matrix functions. Under the assumption that the minimal block size in the divide-and-conquer algorithm is $\mathcal{O}(1)$ and the low-rank update in the Krylov method converges in $\mathcal{O}(1)$ iterations, the algorithm runs with complexity $\mathcal{O}(nk^{2})$ for $k$-banded matrices.
@@
There are other related works that use matrix-vector products to recover matrices with certain properties or using compressed sensing to recover matrices with only few large entries. In [40, 46, 57], the authors devise algorithms or strategies based on matrix-vector products for recovering structured matrices such as hierarchical low-rank matrices and Toeplitz matrices or for answering queries such as whether an unknown matrix is symmetric or not. In particular, Curtis, Powell and Reid [19] and later Coleman and Moré [16] use a certain graph coloring of the adjacency graph of a sparse or banded matrix to approximate sparse or banded Jacobian matrices using matrix-vector products. Bekas, Kokiopoulou and Saad [2] suggested using the same set of coloring as the banded case to recover matrices with rapid off-diagonal decay.
@@
In [41, 53], the authors formulate an algorithm based on compressed sensing techniques to recover sparse matrices with at most $\mathcal{O}(sn/\log^{2}n)$ nonzero entries using $s$ matrix-vector products (See [53, Thm. 6.3]). This method uses a vectorized form of the matrix in question after a certain transformation, which can become expensive. Lastly, Dasarathy, Shah and Bhaskar [20] use compressed sensing techniques by applying compression (sensing operators) on both sides, i.e. $\bm{Y}_{1}\bm{X}\bm{Y}_{2}^{T}$, to recover an unknown sparse matrix $\bm{X}$. The authors use $\mathcal{O}(\sqrt{n}\log n)$ matrix-vector products with $\bm{X}$ where the number of nonzero entries of $\bm{X}$ is $\mathcal{O}(n)$. This can become prohibitive for large $n$ as computing matrix-vector products is usually the dominant cost.
@@ 1.2 Contributions
Our first contribution is devising strategies for approximating (or sometimes recovering exactly, e.g. for sparse matrices) functions of sparse matrices and matrices with similar structures, e.g., matrices with only few large entries in each row/column or matrices with rapid off-diagonal decay. We provide two strategies, one for functions of sparse matrices or matrices with only few large entries in each row/column and another for functions of banded matrices or matrices with off-diagonal decay. The two strategies are based on approximating the large entries of the matrix in question using matrix-vector products only. This is advantageous if the matrix dimension is too large to store explicitly or the matrix is only available through matrix-vector products. For general sparse matrices and their functions, we use compressed sensing to estimate the large entries of each row/column without needing the sparsity pattern of $\bm{A}$. For the banded case, our work is analogous to the previous work done in [16, 19, 31] but we give a focused treatment of banded matrices.
Our second contribution is designing simple algorithms, Algorithm 1 (SpaMRAM) for sparse matrices with general sparsity pattern and Algorithm 2 (BaMRAM) for banded matrices, with the $2$-norm error bounds based on the decay structure of the matrix. We show that SpaMRAM converges quickly as long as the matrix has only few large entries in each row/column and that BaMRAM converges exponentially with the number of matrix-vector products. Assuming that $f$ is sufficiently smooth so that the Krylov method converges in $\mathcal{O}(1)$ iterations when computing $f(\bm{A})\bm{b}$, SpaMRAM runs with complexity $\mathcal{O}(n^{2}\log n)$ for functions of general sparse matrices and BaMRAM runs linearly in $n$ with complexity $\mathcal{O}(nk^{2})$ for functions of $k$-banded matrices. For banded matrices, BaMRAM runs with complexity $\approx 4nk^{2}$ and for functions of banded matrices, BaMRAM has the same complexity, $\mathcal{O}(nk^{2})$, as [18]. The output of the algorithm is also sparse and has $\mathcal{O}(nk)$ nonzero entries where $k$ is the number of nonzero entries recovered per row/column, which can be advantageous in many situations as $f(\bm{A})$ need not be exactly sparse even if $\bm{A}$ is sparse.

### v3

@@ Abstract
[…]
@@ 1 Introduction
The evaluation of matrix functions is an important task arising frequently in many areas of scientific computing. They appear in the numerical solution of partial differential equations pde1; pde2; pde3, network analysis netanal1; EstradaHigham2010, electronic structure calculations elec2; elec1 and statistical learning CortinovisKressner2021; Wengeretal2022, among others. In some problems, the matrix of interest has some underlying structure such as bandedness or sparsity. For example, in graphs with community structure one requires functions of sparse matrices Newman2006 and for electronic structure calculations one computes functions of banded matrices elec2; elec1. In this paper, we consider the problem of approximating $f(\bm{A})$ through matrix-vector products $\bm{x}\mapsto\bm{Ax}$ only, assuming $\bm{A}$ is sparse or banded.
@@
There are many other equivalent ways of defining $f(\bm{A})$, see for example matfunchigham.
@@
When the matrix dimension is large, explicitly computing $f(\bm{A})$ using standard algorithms matcomp; matfunchigham becomes prohibitive in practice, usually requiring $O(n^{3})$ operations. Moreover, the matrix $\bm{A}$ may also be too large to store explicitly. In this scenario, $\bm{A}$ is often only available through matrix-vector products and $f(\bm{A})\bm{b}$  [^1] We note that the matrix-vector product $f(\bm{A})\bm{b}$ is not computed by explicitly forming $f(\bm{A})$ and then multiplying it by $\bm{b}$. Instead, it is approximated using matrix-vector products involving $\bm{A}$ alone, typically through Krylov subspace methods. is used to infer information about $f(\bm{A})$ Guttel2013; GuttelKressnerLund2020. In this paper, we show that the whole matrix $f(\bm{A})$ can still be approximated in a sparse format using $\ll n$ matrix-vector products with $\bm{A}$ if $\bm{A}$ is sparse and $f$ is smooth. We exploit the decay bounds for the entries of $f(\bm{A})$ Benzi2007 to recover its large entries using matrix-vector products; see also Benzi2016. We primarily study functions of sparse matrices with general sparsity pattern and matrices with few large entries in each row/column (see Section 2), but also consider functions of banded matrices and matrices with off-diagonal decay as a special case (see Section 3). Our work is particularly useful if $f(\bm{A})$ is sparse or has only a few entries that are large in each row/column or if $\bm{A}$ is banded with bandwidth $k\ll n$. Throughout this work, we assume that the function $f$ is analytic on the closure of a domain that contains the eigenvalues of $\bm{A}$.
@@
Let $\bm{A}$ be a diagonalizable, sparse matrix with some sparsity pattern and $f$ (either $f:\mathbb{R}\rightarrow\mathbb{R}$ or $f:\mathbb{C}\rightarrow\mathbb{C}$) a function that is analytic in a region containing the spectrum of $\bm{A}$. Depending on the sparsity pattern of $\bm{A}$ and the smoothness of $f$, some entries of $f(\bm{A})$ can be shown to be exponentially smaller than others Benzi2007 and we can view them as noise. For a sparse matrix with a general sparsity pattern, it can be difficult to know the support (location of the large entries) of $f(\bm{A})$ in advance, for example, if the sparsity pattern of $\bm{A}$ is unknown. However, we can write $f(\bm{A})=\widehat{f(\bm{A})}+\bm{E}$ with the matrix $\bm{E}$ having only small entries in magnitude and compute $\widehat{f(\bm{A})}$ by viewing $\widehat{f(\bm{A})}$ as signals and $\bm{E}$ as noise using compressed sensing row by row or column by column. Compressed sensing is a technique in signal processing for efficiently reconstructing a signal which is sparse in some domain by finding solutions to underdetermined linear systems CandesRombergTao06; CandesTao06; Donoho06. As a special case (by taking $f(\bm{A})=\bm{A}$), our algorithm can be used for recovering a sparse matrix $\bm{A}$ itself, where $\bm{A}$ is accessed only through matrix-vector multiplications $\bm{x}\mapsto\bm{A}\bm{x}$. The details will be discussed in Section 2.
@@
Let $\bm{A}$ now be a banded matrix. Then $f(\bm{A})$ can usually be well approximated by a banded matrix DMS1984. There are theoretical results that confirm this property BenziGolub, which show that the entries of $f(\bm{A})$ decay exponentially away from the main diagonal. The observation is that $\left|[f(\bm{A})]_{ij}\right|\ll\left|[f(\bm{A})]_{ii}\right|,\left|[f(\bm{A})]_{jj}\right|$ for $|i-j|\gg 1$, that is, the large entries are typically located not too far away from the main diagonal. Therefore, we can dismiss the exponentially small entries and find a matrix that recovers the large entries of $f(\bm{A})$, i.e., find a matrix $\widehat{f(\bm{A})}$ such that the difference $f(\bm{A})-\widehat{f(\bm{A})}$ only contains small entries in magnitude. We can show that the large entries of $f(\bm{A})$ can be captured through matrix-vector products with the following simple deterministic matrix
@@
where $s<n$, $\bm{I}_{s}$ is the $s\times s$ identity matrix and $\bm{I}_{s}(:,1:\{n/s\}\cdot s)$ are the first $\{n/s\}\cdot s$ columns of $\bm{I}_{s}$, which ensures that $\bm{I}_{n}^{(s)}$ has exactly $n$ rows. Here $\{x\}$ is the fractional part of $x$. This is unlike the sparse case where we now have significant information on the support of $f(\bm{A})$. This set of vectors $\bm{I}_{n}^{(s)}$ are identical to the probing vectors proposed in FrommerSchnimmelSchweitzer2021 for estimating functions of banded matrices and ColemanMore1983; CPR1974 for estimating sparse, banded Jacobian matrices. If we let $s$ be proportional to the bandwidth of $\bm{A}$, the matrix-vector products $f(\bm{A})\bm{I}_{n}^{(s)}$ capture the large entries near the main diagonal, giving us a good approximation for $f(\bm{A})$. We discuss the details in Section 3.
Both cases require matrix-vector products with $f(\bm{A})$, i.e., $\bm{x}\mapsto f(\bm{A})\bm{x}$. Computing this matrix-vector product to high accuracy is crucial for our algorithms. Popular methods include the standard (polynomial) Krylov method matfunchigham; Saad92, and contour integration matveccontour. In this paper, we discuss three functions: the exponential $e^{\bm{A}}$, the square root function $\sqrt{\bm{A}}$ and the logarithm $\log(\bm{A})$. For the exponential we use the polynomial Krylov method, and for the square root function and the logarithm, we use contour integration as in (matveccontour, method 2) as both functions have a branch cut on the negative real axis $(-\infty,0]$, making them difficult to approximate using polynomials if $\bm{A}$ has eigenvalues close to 0. For Krylov methods, block versions can also be used FrommerLundSzyld2018, which compute $f(\bm{A})\bm{B}\in\mathbb{R}^{n\times s}$ at once instead of column by column. Although the computation of the matrix-vector product $f(\bm{A})\bm{b}$ is important and crucial to the approximation of $f(\bm{A})$, we do not discuss the details further as they are not the focus of this paper; see for example (matfunchigham, Ch.13).
@@ 1.1 Existing methods
First, Benzi and Razouk Benzi2007 describe an algorithm that exploits the decay property by first finding a good polynomial approximation $p$ to $f$, for example using Chebyshev interpolation, and computing $p(\bm{A})$ while using a dropping strategy to keep the matrix as sparse as possible. However, this method can quickly become infeasible if $f$ cannot be well approximated by a low-degree polynomial or if sparse data structures and arithmetic is not available. Frommer, Schimmel and Schweitzer FrommerSchnimmelSchweitzer2021 use probing methods to approximate matrix functions $f(\bm{A})$ and its related quantities, for instance the trace of $f(\bm{A})$. The method finds probing vectors to estimate the entries of $f(\bm{A})$ by partitioning the vertices of the directed graph $G(\bm{A})$ associated with the sparse matrix $\bm{A}$ using graph coloring and the sparsity pattern of $\bm{A}$. Our Section 3 on banded matrices can be seen as a special case of FrommerSchnimmelSchweitzer2021, but as banded matrices arise naturally in applications and come with strong theoretical results, we make the sensing matrix explicit and study them in detail. The strategy proposed in FrommerSchnimmelSchweitzer2021 costs $\mathcal{O}(\Delta^{d}n)$ to obtain the graph coloring where $\Delta$ is the maximal degree of $G(\bm{A})$ and $d$ is a distance parameter. In the special case when $\bm{A}$ is a banded matrix the graph coloring can be obtained with $\mathcal{O}(n)$ cost. Our work and FrommerSchnimmelSchweitzer2021 satisfies a similar theoretical accuracy bound up to a low-degree polynomial factor in $n$ (cf. Section 3 of FrommerSchnimmelSchweitzer2021 and Theorem 1 in our work), but we use compressed sensing to estimate the large entries of each row/column. Cortinovis, Kressner and Massei CortinovisKressnerMassei2022 use a divide-and-conquer approach to approximate matrix functions where the matrix is either banded, hierarchically semiseparable or has a related structure. The divide-and-conquer method is based on (rational) Krylov subspace methods for performing low-rank updates of matrix functions. Under the assumption that the minimal block size in the divide-and-conquer algorithm is $\mathcal{O}(1)$ and the low-rank update in the Krylov method converges in $\mathcal{O}(1)$ iterations, the algorithm runs with complexity $\mathcal{O}(nk^{2})$ for $k$-banded matrices.
@@
There are other related works that use matrix-vector products to recover matrices with certain properties or using compressed sensing to recover matrices with only few large entries. In Townsend22; LetvittMartinsson2022; Woodruffetal2021, the authors devise algorithms or strategies based on matrix-vector products for recovering structured matrices such as hierarchical low-rank matrices and Toeplitz matrices or for answering queries such as whether an unknown matrix is symmetric or not. In particular, Curtis, Powell and Reid CPR1974 and later Coleman and Moré ColemanMore1983 use a certain graph coloring of the adjacency graph of a sparse or banded matrix to approximate sparse or banded Jacobian matrices using matrix-vector products. Bekas, Kokiopoulou and Saad BekasKokiopoulouSaad2007 suggested using the same set of coloring as the banded case to recover matrices with rapid off-diagonal decay.
@@
In HermanStrohmer2009; PfranderRauhutTanner2008, the authors formulate an algorithm based on compressed sensing techniques to recover sparse matrices with at most $\mathcal{O}(sn/\log^{2}n)$ nonzero entries using $s$ matrix-vector products (See (PfranderRauhutTanner2008, Thm. 6.3)). This method uses a vectorized form of the matrix in question after a certain transformation, which can become expensive. Lastly, Dasarathy, Shah and Bhaskar Dasarathyetal2015 use compressed sensing techniques by applying compression (sensing operators) on both sides, i.e. $\bm{Y}_{1}\bm{X}\bm{Y}_{2}^{T}$, to recover an unknown sparse matrix $\bm{X}$. The authors use $\mathcal{O}(\sqrt{n}\log n)$ matrix-vector products with $\bm{X}$ where the number of nonzero entries of $\bm{X}$ is $\mathcal{O}(n)$. This can become prohibitive for large $n$ as computing matrix-vector products is usually the dominant cost.
@@ 1.2 Contributions
Our first contribution is devising strategies for approximating (or sometimes recovering exactly, e.g. for sparse matrices) functions of sparse matrices and matrices with similar structures, e.g., matrices with only few large entries in each row/column or matrices with rapid off-diagonal decay, which is typically the case when the function $f$ is analytic on the closure of a domain that contains the eigenvalues of $\bm{A}$. We provide two strategies, one for functions of sparse matrices or matrices with only few large entries in each row/column and another for functions of banded matrices or matrices with off-diagonal decay. The two strategies are based on approximating the large entries of the matrix in question using matrix-vector products only. This is advantageous if the matrix dimension is too large to store explicitly or the matrix is only available through matrix-vector products. For general sparse matrices and their functions, we use compressed sensing to estimate the large entries of each row/column without needing the sparsity pattern of $\bm{A}$. For the banded case, our work is analogous to the previous work done in ColemanMore1983; CPR1974; FrommerSchnimmelSchweitzer2021 but we give a focused treatment of banded matrices.
Our second contribution is designing simple algorithms, Algorithm 1 (SpaMRAM) for sparse matrices with general sparsity pattern and Algorithm 2 (BaMRAM) for banded matrices, with the $2$-norm error bounds based on the decay structure of the matrix. We show that SpaMRAM converges quickly as long as the matrix has only few large entries in each row/column and that BaMRAM converges exponentially with the number of matrix-vector products. Assuming that $f$ is sufficiently smooth so that the Krylov method converges in $\mathcal{O}(1)$ iterations when computing $f(\bm{A})\bm{b}$, SpaMRAM runs with complexity $\mathcal{O}(n^{2}\log n)$ for functions of general sparse matrices and BaMRAM runs linearly in $n$ with complexity $\mathcal{O}(nk^{2})$ for functions of $k$-banded matrices. For banded matrices, BaMRAM runs with complexity $\approx 4nk^{2}$ and for functions of banded matrices, BaMRAM has the same complexity, $\mathcal{O}(nk^{2})$, as CortinovisKressnerMassei2022. The output of the algorithm is also sparse and has $\mathcal{O}(nk)$ nonzero entries where $k$ is the number of nonzero entries recovered per row/column, which can be advantageous in many situations as $f(\bm{A})$ need not be exactly sparse even if $\bm{A}$ is sparse.